# A globalized inexact semismooth Newton method for strongly convex optimal control problems

### v3

@@
Abstract. We investigate a globalized inexact semismooth Newton method applied to strongly convex optimization problems in Hilbert spaces. Here, the semismooth Newton method is appplied to the dual problem, which has a continuously differentiable objective. We prove global strong convergence of iterates as well as transition to local superlinear convergence. The latter needs a second-order Taylor expansion involving semismooth derivative concepts. The convergence of the globalized method is demonstrated in numerical examples, for which the local unglobalized method diverges.
@@ 1 Introduction
$\min_{u\in U}\frac{1}{2}\|Su-z\|_{Y}^{2}+\frac{\alpha}{2}\|u\|_{U}^{2}+g(u)$ (1.1)
@@
Following [18, 42, 25], we propose to apply the semismooth Newton method to the dual problem of (1.1), where the globalization uses the dual objective as merit function. Due to the structure of (1.1), the dual problem has a continuously differentiable objective functional, which enables a particularly nice convergence analysis. As this analysis crucially relies on the strong convexity of the primal problem, it cannot be extended to non-convex or not strongly convex problems. Hence, the proposed method is intended to be used as a subproblem solver in, e.g., proximal point methods [25] or iterated Bregman methods [28].
@@
$u=\operatorname{prox}_{g/\alpha}\left(-\frac{1}{\alpha}S^{*}(Su-z)\right),$ (1.2)
@@
$\xi-z+S\operatorname{prox}_{g/\alpha}(S^{*}(\xi/\alpha))=0,$ (1.3)
@@
A third approach to solve (1.1) is based on the so-called normal map [33], where one solves
$p-S^{*}(S\operatorname{prox}_{g/\alpha}(-p/\alpha)-z)=0$ (1.4)
@@
This artice is written 25 years after the publication of the seminal contributions [39] (submitted 2000) and [14] (submitted 2001) on semismooth Newton methods applied to optimal control problems. Still the literature on globalized semismooth Newton methods for infinite-dimensional problems is rather scarce. In fact, to the best of our knowledge, only the following contributions develop a global convergent method applicable to (1.1). In [18] the globalization based on (1.3) was investigated for the special case, where $g$ is the indicator function associated with box constraints. In the monograph [40] a trust-region globalization of the semi-smooth Newton method applied to the non-smooth equation (1.2) is analyzed. Since continuous differentiability of $g$ is required, it is not applicable to choices like $g(u)=\|u\|_{L^{1}(\Omega)}$. A similar method was analyzed for the special case of a parabolic control problem in [20]. In [8] a globalized method is introduced, where the convergence analysis relies on strong assumptions on the sequence of chosen derivatives. For $\ell^{1}$-regularized problems, a globalized semismooth Newton method can be found in [12]. One major drawback of the convergence analysis in [40, 8, 12] is that they assume the existence of strong accumulation points of the sequence of iterates. While it is natural to expect that the sequence of iterates is bounded, this only guarantees the existence of weakly converging subsequences. However, strong convergence of iterates has to be verified in order to prove transition to fast local convergence.
@@
Let us comment on related work on globalized semismooth Newton methods. The recent work [1] introduces a semismooth Newton method to solve contractive fixed point problems. Due to the contraction assumption, no damping or globalization is needed. In [30, 29] globalized proximal Newton methods that solve minimization problems in Hilbert spaces are investigated. The use of the normal map equation was proposed in [17] together with a damped Newton method without convergence analysis. Global convergence of the unregularized semismooth Newton method applied to problems with box constraints was proven in [3, 14] for discretized problems and [22] in the infinite-dimensional setting provided $\alpha$ is large relative to $\|S\|$. For a review of early results on (globalized) semismooth Newton methods we refer to [31]. Recent contributions on globalized semismooth Newton methods to minimize the sum of a differentiable and a non-smooth function in finite dimensions can be found in [26, 9].
@@
The goal of this article is to give an almost self-contained analysis of the global convergence of the globalized semismooth Newton method. A few facts about subdifferentials and proximal operators are collected in Section 2.1. The dual problem to (1.1) is introduced in Section 2.2. Due to the problem structure, no duals have to be computed for the implementation, only the proximal operator of $g$ is needed. For the transition to fast convergence, a second order Taylor expansion is essential, which is introduced in Section 3. As in [42, 25] we use a conjugate gradient method as solver for the Newton equations. In order to prove that all iterates of the CG method are descent directions, one elementary result is needed, which is provided in Section 4. The main section is Section 5, which contains the globalized semismooth Newton method Algorithm 2. We prove strong convergence of the iterates in Lemma 5.10, and the transition to fast local convergence in Theorem 5.13. Section 6 shows that commonly studied examples admit a second order Taylor expansion. As we show in Section 7, the application of the semismooth Newton method to (1.3) enables the use of the variational discretization [15] similar to the arguments in [17]. The paper concludes with numerical experiments in Section 8. The results indicate that the variational discretization leads to better performance of the semismooth Newton method for small values of $\alpha$ in some examples.
@@ 2 Preliminaries on the primal and dual problems
$0\in S^{*}(S\bar{u}-\color[rgb]{1,0,0}z\color[rgb]{0,0,0})+\alpha\bar{u}+\partial g(\bar{u}).$ (2.1)
Proof.
Existence can be proven by classical arguments, see, e.g., [2, Theorem 11.10]. Let $\bar{u}$ be a solution of (1.1). Since the first two addends of (1.1) are continuous, we can apply the subgradient sum-rule twice to obtain (2.1). The sufficiency of (2.1) follows from convexity. ∎
@@ 2.1 Conjugate functions and proximal operators
For properties of Moreau-Yosida envelopes and proximal operators, we refer to [2, Section 12.4]. Convex conjugates are developed in [2, Chapter 13].
Proposition 2.3 ([2, Proposition 12.26]).
@@
Due to Proposition 2.5, it follows that $H^{*}$ is Fréchet differentiable, and the expression of $\nabla H^{*}$ follows. Boundedness from below is a direct consequence of the definition of $H^{*}$. ∎
@@ 2.2 Optimality conditions for the dual problem
Let us introduce the Fenchel dual problem to (1.1), see, e.g., [2, Section 15.3]. The dual problem of (1.1) is given by
$\min_{\xi\in Y}F^{*}(-\xi)+G^{*}(S^{*}\xi)=:\Phi(\xi)$ (2.2)
@@
Proof.
The function $\Phi$ is continuous and convex. Since $F^{*}$ is coercive, existence of solutions can be proven easily. Uniqueness of minimizers is a consequence of the strict convexity of $F^{*}$. The dual problem (2.2) is the minimization problem of the convex and Fréchet differentiable function $\Phi$. Then $\nabla\Phi(\bar{\xi})=0$ is necessary and sufficient for optimality of $\bar{\xi}$. The expression of the derivative is a consequence of Proposition 2.5. ∎
@@
Proof.
Due to Proposition 2.3, the optimality condition (2.1) is equivalent to
@@
Then the claim follows by elementary computations. ∎
@@ 3 Second-order semismoothness
For the proof of transition from global convergence to fast local convergence, a second-order Taylor expansion of the function $G^{*}$ is indispensible [7]. Following [30], we introduce the notion of second-order semismoothness.
@@
$\sup_{M\in\partial^{2}T(x+h)}\left|T(x+h)-T(x)-T^{\prime}(x)h-\frac{1}{2}M(h,h)\right|=o(\|h\|_{X}^{2})$ (3.1)
@@
$\sup_{M\in\partial^{2}T(x+h)}\left\|T^{\prime}(x+h)-T^{\prime}(x)-M(h,\cdot)\right\|_{X^{*}}=o(\|h\|_{X})$ (3.2)
@@
Here, second-order semismoothness refers to the availability of a second-order Taylor expansion (3.1), which is different to the use in [40], where higher-order semismoothness refers to stronger estimates of the first-order expansion (3.2) replacing $o(\|h\|_{X})$ for $O(\|h\|_{X}^{1+\alpha})$ with $\alpha\in(0,1]$.
If $X$ is a Hilbert space, we can use operators $M\in\mathcal{L}(X,X)$ instead of bilinear forms, and replace $M(h,h)$ and $M(h,\cdot)$ by $\langle Mh,h\rangle$ and $Mh$. In [30] an example is presented to show that (3.1) does not imply (3.2). Let us prove a sufficient condition for (3.1) which is only based on assumptions on $T^{\prime}$.
@@
$T(x+h)-T(x)-\frac{1}{2}[T^{\prime}(x)+T^{\prime}(x+h)]h\\ \begin{aligned} &=\int_{0}^{1}[T^{\prime}(x+sh)-T^{\prime}(x)]h\,\text{{}{}d}s-\frac{1}{2}[T^{\prime}(x+h)-T^{\prime}(x)]h\\ &=\int_{0}^{1}[sT^{\prime\prime}(x;h)+R_{s}]h\,\text{{}{}d}s-\frac{1}{2}T^{\prime\prime}(x;h)-\frac{1}{2}R_{1}h\\ &=\int_{0}^{1}R_{s}h\,\text{{}{}d}s-\frac{1}{2}R_{1}h\end{aligned}$
@@
which finishes the proof. ∎
@@ 4 Conjugate Gradients
Let us recall the celebrated conjugate gradient method to solve linear systems with positive definite operators. Let $X$ be a Hilbert space, $A\in\mathcal{L}(X,X)$ self-adjoint and positive definite, and $b\in X$. The CG algorithm to solve $Ax=b$ in the notation from [38, Section 38] can be found in Algorithm 1.
@@
The classical orthogonality relations and convergence guarantees transfer to the Hilbert space case, see, e.g., [6]. If $A=I+K$ with $K$ compact we have r-superlinear convergence of $x_{k}\to A^{-1}b$ [41]. These results imply that for positive tolerance $\delta_{\mathrm{CG}}$, algorithm Algorithm 1 terminates after finitely many steps.
In the following, we need an estimate of $\langle x_{k},b\rangle$. A similar result can be found in [42, Lemma 3.1]. We give a simplified proof.
@@
Let $A\in\mathcal{L}(X,X)$ be self-adjoint and positive definite, and $b\in X$. Let $(x_{k})$ be iterates of Algorithm 1. Then
@@
Due to the orthogonality relations of the residuals, [38, (38.4)], we have $\langle r_{k},b\rangle=\langle r_{k},r_{0}\rangle=0$ for all $k\geq 1$. Then we get for the conjugate directions
@@
so that the sequence $(\langle x_{k},b\rangle)$ is monotonically increasing. ∎
@@ 5 Semismooth Newton
$\Phi(\xi)=\frac{1}{2}\|\xi-z\|_{Y}^{2}-\frac{1}{2}\|z\|_{Y}^{2}+\frac{1}{2\alpha}\|S^{*}\xi\|_{U}^{2}-\alpha\operatorname{env}_{g/\alpha}(S^{*}\xi/\alpha).$ (5.1)
Its gradient is given by Lemma 2.7 as
$\nabla\Phi(\xi)=\xi-z+S\operatorname{prox}_{g/\alpha}(S^{*}\xi/\alpha).$ (5.2)
@@
Proof.
Differentiability of $\Phi$ is a consequence of the differentiability of $\operatorname{env}_{g/\alpha}$ by Proposition 2.5. By Proposition 2.4, the proximal operator $\operatorname{prox}_{g/\alpha}$ is Lipschitz continuous. Hence, $\nabla\Phi$ is Lipschitz continuous. The claimed estimate follows is then a consequence of this Lipschitz property. ∎
@@
$\Psi(\xi):=\frac{1}{2\alpha}\|S^{*}\xi\|_{U}^{2}-\alpha\operatorname{env}_{g/\alpha}(S^{*}\xi/\alpha),$ (5.3)
@@
such that $\Psi$ given by (5.3) is second-order semismooth with respect to $\partial^{2}\Psi$, see Definition 3.1. In addition, $\partial^{2}\Psi$ is bounded, i.e., there is $c>0$ such that for all $y\in Y$ and $M\in\partial^{2}\Psi(y)$ it holds
@@
Let us emphasize that in the algorithm we will not need to know the precise values of the Lipschitz constant $L$ or the bound $c$. For the particular case of Nemyzki operators on $L^{2}$ spaces, validity of 1 is discussed in Section 6.
@@
$(I+M_{k})d_{k}=-\nabla\Phi(\xi_{k})=-\left(\xi_{k}-z+S\operatorname{prox}_{g/\alpha}(S^{*}\xi_{k}/\alpha)\right).$ (5.4)
In the upcoming convergence analysis, we will use both remainder estimates (3.1) and (3.2). The first-order estimate is an ingredient of the classical local convergence proof of semismooth Newton [40], while the second-order estimate is needed in the proof of transition to fast local convergence with full steps ($t_{k}=1$).
@@
  Choose $M_{k}\in\partial^{2}\Phi(\xi_{k})$, see Corollary 5.2.
@@
$\|M_{k}d_{k}+\nabla\Phi(\xi_{k})\|_{Y}\leq\eta\|\nabla\Phi(\xi_{k})\|_{Y}^{1+\tau}$ (5.5)
  using Algorithm 1 with tolerance $\triangleright$ CG
@@
$\Phi(\xi_{k}+\beta^{l_{k}}d_{k})-\Phi(\xi_{k})\leq\sigma\beta^{l_{k}}\langle d_{k},\nabla\Phi(\xi_{k})\rangle$ (5.6)
@@
Algorithm 2 is basically [25, Algorithm 4.1]. The convergence proof in the finite-dimensional case is given in [25, Theorem 4.1]. For the special case of box constraints and exact Newton solves ($\eta=0$), Algorithm 2 is equivalent to [18, Algorithm 4.2].
@@
Algorithm 2 is well-defined.
Proof.
Suppose $\nabla\Phi(\xi_{k})\neq 0$. Due to the special form of the Newton derivatives $M_{k}\in\partial^{2}\Phi(\xi_{k})$ in Corollary 5.2, it follows directly that the operators $M_{k}$ are self-adjoint, positive definite, and invertible. The direction $M_{k}^{-1}\nabla\Phi(\xi_{k})$, satisfies (5.5). Then Algorithm 1 terminates after finitely many steps. Due to Lemma 4.1 the returned element $d_{k}$ is a descent direction to the Fréchet differentiable function $\Phi$. Under this condition, the Armijo linesearch terminates after finitely many steps. ∎
The termination criterion is motivated by the following bound, see also [18, Lemma 4.3 (iii)].
@@
which proves the claim using Cauchy-Schwarz inequality. ∎
@@
Let $(\xi_{k})$ be a sequence of iterates of Algorithm 2 together with the search directions $(d_{k})$ satisfying (5.5). Then it holds
$\langle d_{k},\nabla\Phi(\xi_{k})\rangle\leq-\frac{1}{1+c}\|\nabla\Phi(\xi_{k})\|_{Y}^{2},$ (5.7)
where $c$ is from 1.
Proof.
Lemma 4.1 implies $\langle d_{k},-\nabla\Phi(\xi_{k})\rangle\geq\frac{1}{\|M_{k}\|_{\mathcal{L}(Y,Y)}}\|\nabla\Phi(\xi_{k})\|_{Y}^{2}$, and the claim follows with Corollary 5.3. ∎
@@
Let $(\xi_{k})$ be a sequence of iterates of Algorithm 2 together with the step sizes $(t_{k})$ and search directions $(d_{k})$. Then
@@
Proof.
Due to Lemma 2.6, $G^{*}$, $\Psi$, and $\Phi$ are bounded from below, hence $(\Phi(\xi_{k}))$ is monotonically decreasing and converging. The claim is now a consequence of (5.6). ∎
@@
Using Corollary 5.1, we will prove a lower bound of $t_{k}$ next. There, we will use the global Lipschitz continuity of $\nabla\Phi$. If we would know that $(\xi_{k})$ has strongly converging subsequences then local Lipschitz continuity of $\nabla\Phi$ would be enough.
@@
Let $(\xi_{k})$ be a sequence of iterates of Algorithm 2 together with the step sizes $(t_{k})$ and search directions $(d_{k})$. Then there is a constant $\sigma^{\prime}>0$ such that $t_{k}<1$ implies
@@
If $t_{k}<1$ satisfies the Armijo condition (5.6) then $l_{k}\geq 1$, and $\beta^{l_{k}-1}=t_{k}/\beta$ violates the Armijo condition (5.6). Consequently, $t_{k}/\beta$ is larger than the quantity on the right-hand side in the above inequality, which implies the claim. ∎

### v4

@@
Abstract
We investigate a globalized inexact semismooth Newton method applied to strongly convex optimization problems in Hilbert spaces. Here, the semismooth Newton method is appplied to the dual problem, which has a continuously differentiable objective. We prove global strong convergence of iterates as well as transition to local superlinear convergence. The latter needs a second-order Taylor expansion involving semismooth derivative concepts. The convergence of the globalized method is demonstrated in numerical examples, for which the local unglobalized method diverges.
@@ 1 Introduction
(1.1) $\min_{u\in U}\frac{1}{2}\|Su-z\|_{Y}^{2}+\frac{\alpha}{2}\|u\|_{U}^{2}+g(u)$
@@
Following [HinzeVierling2012oms, ZhaoSunToh2010, MilzarekSchaippUlbrich2024], we propose to apply the semismooth Newton method to the dual problem of (1.1), where the globalization uses the dual objective as merit function. Due to the structure of (1.1), the dual problem has a continuously differentiable objective functional, which enables a particularly nice convergence analysis. As this analysis crucially relies on the strong convexity of the primal problem, it cannot be extended to non-convex or not strongly convex problems. Hence, the proposed method is intended to be used as a subproblem solver in, e.g., proximal point methods [MilzarekSchaippUlbrich2024] or iterated Bregman methods [PornerWachsmuth2016].
@@
(1.2) $u=\operatorname{prox}_{g/\alpha}\left(-\frac{1}{\alpha}S^{*}(Su-z)\right),$
@@
(1.3) $\xi-z+S\operatorname{prox}_{g/\alpha}(S^{*}(\xi/\alpha))=0,$
@@
A third approach to solve (1.1) is based on the so-called normal map [Robinson1992], where one solves
(1.4) $p-S^{*}(S\operatorname{prox}_{g/\alpha}(-p/\alpha)-z)=0$
@@
This artice is written 25 years after the publication of the seminal contributions [Ulbrich2002] (submitted 2000) and [HintermullerItoKunisch2002] (submitted 2001) on semismooth Newton methods applied to optimal control problems. Still the literature on globalized semismooth Newton methods for infinite-dimensional problems is rather scarce. In fact, to the best of our knowledge, only the following contributions develop a global convergent method applicable to (1.1). In [HinzeVierling2012oms] the globalization based on (1.3) was investigated for the special case, where $g$ is the indicator function associated with box constraints. In the monograph [Ulbrich2011] a trust-region globalization of the semi-smooth Newton method applied to the non-smooth equation (1.2) is analyzed. Since continuous differentiability of $g$ is required, it is not applicable to choices like $g(u)=\|u\|_{L^{1}(\Omega)}$. A similar method was analyzed for the special case of a parabolic control problem in [KelleySachs1999]. In [GerdtsHornKimmerle2017] a globalized method is introduced, where the convergence analysis relies on strong assumptions on the sequence of chosen derivatives. For $\ell^{1}$-regularized problems, a globalized semismooth Newton method can be found in [HansRaasch2015]. One major drawback of the convergence analysis in [Ulbrich2011, GerdtsHornKimmerle2017, HansRaasch2015] is that they assume the existence of strong accumulation points of the sequence of iterates. While it is natural to expect that the sequence of iterates is bounded, this only guarantees the existence of weakly converging subsequences. However, strong convergence of iterates has to be verified in order to prove transition to fast local convergence.
@@
Let us comment on related work on globalized semismooth Newton methods. The recent work [alphonse2024globalizedinexactsemismoothnewton] introduces a semismooth Newton method to solve contractive fixed point problems. Due to the contraction assumption, no damping or globalization is needed. In [PotzlSchielaJaap2022, PotzlSchielaJaap2024] globalized proximal Newton methods that solve minimization problems in Hilbert spaces are investigated. The use of the normal map equation was proposed in [HinzeVierling_2012spp] together with a damped Newton method without convergence analysis. Global convergence of the unregularized semismooth Newton method applied to problems with box constraints was proven in [BergouniouxItoKunisch1999, HintermullerItoKunisch2002] for discretized problems and [KunischRosch2002] in the infinite-dimensional setting provided $\alpha$ is large relative to $\|S\|$. For a review of early results on (globalized) semismooth Newton methods we refer to [QiSun1999]. Recent contributions on globalized semismooth Newton methods to minimize the sum of a differentiable and a non-smooth function in finite dimensions can be found in [OuyangMilzarek2024, Gfrerer2025].
@@
The goal of this article is to give an almost self-contained analysis of the global convergence of the globalized semismooth Newton method. A few facts about subdifferentials and proximal operators are collected in Section˜2.1. The dual problem to (1.1) is introduced in Section˜2.2. Due to the problem structure, no duals have to be computed for the implementation, only the proximal operator of $g$ is needed. For the transition to fast convergence, a second order Taylor expansion is essential, which is introduced in Section˜3. As in [ZhaoSunToh2010, MilzarekSchaippUlbrich2024] we use a conjugate gradient method as solver for the Newton equations. In order to prove that all iterates of the CG method are descent directions, one elementary result is needed, which is provided in Section˜4. The main section is Section˜5, which contains the globalized semismooth Newton method Algorithm˜2. We prove strong convergence of the iterates in Lemma˜5.17, and the transition to fast local convergence in Theorem˜5.23. Section˜6 shows that commonly studied examples admit a second order Taylor expansion. As we show in Section˜7, the application of the semismooth Newton method to (1.3) enables the use of the variational discretization [Hinze2005] similar to the arguments in [HinzeVierling_2012spp]. The paper concludes with numerical experiments in Section˜8. The results indicate that the variational discretization leads to better performance of the semismooth Newton method for small values of $\alpha$ in some examples.
@@ 2 Preliminaries on the primal and dual problems
(2.1) $0\in S^{*}(S\bar{u}-z)+\alpha\bar{u}+\partial g(\bar{u}).$
Proof 2.2.
Existence can be proven by classical arguments, see, e.g., [BauschkeCombettes2017, Theorem 11.10]. Let $\bar{u}$ be a solution of (1.1). Since the first two addends of (1.1) are continuous, we can apply the subgradient sum-rule twice to obtain (2.1). The sufficiency of (2.1) follows from convexity.
@@ 2.1 Conjugate functions and proximal operators
For properties of Moreau-Yosida envelopes and proximal operators, we refer to [BauschkeCombettes2017, Section 12.4]. Convex conjugates are developed in [BauschkeCombettes2017, Chapter 13].
Proposition 2.4 ([BauschkeCombettes2017, Proposition 12.26]).
@@
Due to Proposition˜2.6, it follows that $H^{*}$ is Fréchet differentiable, and the expression of $\nabla H^{*}$ follows. Boundedness from below is a direct consequence of the definition of $H^{*}$.
@@ 2.2 Optimality conditions for the dual problem
Let us introduce the Fenchel dual problem to (1.1), see, e.g., [BauschkeCombettes2017, Section 15.3]. The dual problem of (1.1) is given by
(2.2) $\min_{\xi\in Y}F^{*}(-\xi)+G^{*}(S^{*}\xi)=:\Phi(\xi)$
@@
Proof 2.10.
The function $\Phi$ is continuous and convex. Since $F^{*}$ is coercive, existence of solutions can be proven easily. Uniqueness of minimizers is a consequence of the strict convexity of $F^{*}$. The dual problem (2.2) is the minimization problem of the convex and Fréchet differentiable function $\Phi$. Then $\nabla\Phi(\bar{\xi})=0$ is necessary and sufficient for optimality of $\bar{\xi}$. The expression of the derivative is a consequence of Proposition˜2.6.
@@
Proof 2.12.
Due to Proposition˜2.4, the optimality condition (2.1) is equivalent to
@@
Then the claim follows by elementary computations.
@@ 3 Second-order semismoothness
For the proof of transition from global convergence to fast local convergence, a second-order Taylor expansion of the function $G^{*}$ is indispensible [Facchinei1995]. Following [PotzlSchielaJaap2022], we introduce the notion of second-order semismoothness.
@@
(3.1) $\sup_{M\in\partial^{2}T(x+h)}\left|T(x+h)-T(x)-T^{\prime}(x)h-\frac{1}{2}M(h,h)\right|=o(\|h\|_{X}^{2})$
@@
(3.2) $\sup_{M\in\partial^{2}T(x+h)}\left\|T^{\prime}(x+h)-T^{\prime}(x)-M(h,\cdot)\right\|_{X^{*}}=o(\|h\|_{X})$
@@
Here, second-order semismoothness refers to the availability of a second-order Taylor expansion (3.1), which is different to the use in [Ulbrich2011], where higher-order semismoothness refers to stronger estimates of the first-order expansion (3.2) replacing $o(\|h\|_{X})$ for $O(\|h\|_{X}^{1+\alpha})$ with $\alpha\in(0,1]$.
If $X$ is a Hilbert space, we can use operators $M\in\mathcal{L}(X,X)$ instead of bilinear forms, and replace $M(h,h)$ and $M(h,\cdot)$ by $\langle Mh,h\rangle$ and $Mh$. In [PotzlSchielaJaap2022] an example is presented to show that (3.1) does not imply (3.2). Let us prove a sufficient condition for (3.1) which is only based on assumptions on $T^{\prime}$.
@@
$T(x+h)-T(x)-\frac{1}{2}[T^{\prime}(x)+T^{\prime}(x+h)]h$ $=\int_{0}^{1}[T^{\prime}(x+sh)-T^{\prime}(x)]h\,\text{{}{}d}s-\frac{1}{2}[T^{\prime}(x+h)-T^{\prime}(x)]h$
$=\int_{0}^{1}[sT^{\prime\prime}(x;h)+R_{s}]h\,\text{{}{}d}s-\frac{1}{2}T^{\prime\prime}(x;h)-\frac{1}{2}R_{1}h$
$=\int_{0}^{1}R_{s}h\,\text{{}{}d}s-\frac{1}{2}R_{1}h$
@@
which finishes the proof.
@@ 4 Conjugate Gradients
Let us recall the celebrated conjugate gradient method to solve linear systems with positive definite operators. Let $X$ be a Hilbert space, $A\in\mathcal{L}(X,X)$ self-adjoint and positive definite, and $b\in X$. The CG algorithm to solve $Ax=b$ in the notation from [TrefethenBau1997, Section 38] can be found in Algorithm˜1.
@@
The classical orthogonality relations and convergence guarantees transfer to the Hilbert space case, see, e.g., [Daniel1967]. If $A=I+K$ with $K$ compact we have r-superlinear convergence of $x_{k}\to A^{-1}b$ [Winther1980]. These results imply that for positive tolerance $\delta_{\mathrm{CG}}$, algorithm Algorithm˜1 terminates after finitely many steps.
In the following, we need an estimate of $\langle x_{k},b\rangle$. A similar result can be found in [ZhaoSunToh2010, Lemma 3.1]. We give a simplified proof.
@@
Let $A\in\mathcal{L}(X,X)$ be self-adjoint and positive definite, and $b\in X$. Let $(x_{k})$ be iterates of Algorithm˜1. Then
@@
Due to the orthogonality relations of the residuals, [TrefethenBau1997, (38.4)], we have $\langle r_{k},b\rangle=\langle r_{k},r_{0}\rangle=0$ for all $k\geq 1$. Then we get for the conjugate directions
@@
so that the sequence $(\langle x_{k},b\rangle)$ is monotonically increasing.
@@ 5 Semismooth Newton
(5.1) $\Phi(\xi)=\frac{1}{2}\|\xi-z\|_{Y}^{2}-\frac{1}{2}\|z\|_{Y}^{2}+\frac{1}{2\alpha}\|S^{*}\xi\|_{U}^{2}-\alpha\operatorname{env}_{g/\alpha}(S^{*}\xi/\alpha).$
Its gradient is given by Lemma˜2.9 as
(5.2) $\nabla\Phi(\xi)=\xi-z+S\operatorname{prox}_{g/\alpha}(S^{*}\xi/\alpha).$
@@
Proof 5.2.
Differentiability of $\Phi$ is a consequence of the differentiability of $\operatorname{env}_{g/\alpha}$ by Proposition˜2.6. By Proposition˜2.5, the proximal operator $\operatorname{prox}_{g/\alpha}$ is Lipschitz continuous. Hence, $\nabla\Phi$ is Lipschitz continuous. The claimed estimate follows is then a consequence of this Lipschitz property.
@@
(5.3) $\Psi(\xi):=\frac{1}{2\alpha}\|S^{*}\xi\|_{U}^{2}-\alpha\operatorname{env}_{g/\alpha}(S^{*}\xi/\alpha),$
@@
such that $\Psi$ given by (5.3) is second-order semismooth with respect to $\partial^{2}\Psi$, see Definition˜3.1. In addition, $\partial^{2}\Psi$ is bounded, i.e., there is $c>0$ such that for all $y\in Y$ and $M\in\partial^{2}\Psi(y)$ it holds
@@
Let us emphasize that in the algorithm we will not need to know the precise values of the Lipschitz constant $L$ or the bound $c$. For the particular case of Nemyzki operators on $L^{2}$ spaces, validity of ˜1 is discussed in Section˜6.
@@
(5.4) $(I+M_{k})d_{k}=-\nabla\Phi(\xi_{k})=-\left(\xi_{k}-z+S\operatorname{prox}_{g/\alpha}(S^{*}\xi_{k}/\alpha)\right).$
In the upcoming convergence analysis, we will use both remainder estimates (3.1) and (3.2). The first-order estimate is an ingredient of the classical local convergence proof of semismooth Newton [Ulbrich2011], while the second-order estimate is needed in the proof of transition to fast local convergence with full steps ($t_{k}=1$).
@@
  Choose $M_{k}\in\partial^{2}\Phi(\xi_{k})$, see Corollary˜5.3.
@@
(5.5) $\|M_{k}d_{k}+\nabla\Phi(\xi_{k})\|_{Y}\leq\eta\|\nabla\Phi(\xi_{k})\|_{Y}^{1+\tau}$
  using Algorithm˜1 with tolerance $\triangleright$ CG
@@
(5.6) $\Phi(\xi_{k}+\beta^{l_{k}}d_{k})-\Phi(\xi_{k})\leq\sigma\beta^{l_{k}}\langle d_{k},\nabla\Phi(\xi_{k})\rangle$
@@
Algorithm˜2 is basically [MilzarekSchaippUlbrich2024, Algorithm 4.1]. The convergence proof in the finite-dimensional case is given in [MilzarekSchaippUlbrich2024, Theorem 4.1]. For the special case of box constraints and exact Newton solves ($\eta=0$), Algorithm˜2 is equivalent to [HinzeVierling2012oms, Algorithm 4.2].
@@
Algorithm˜2 is well-defined.
Proof 5.6.
Suppose $\nabla\Phi(\xi_{k})\neq 0$. Due to the special form of the Newton derivatives $M_{k}\in\partial^{2}\Phi(\xi_{k})$ in Corollary˜5.3, it follows directly that the operators $M_{k}$ are self-adjoint, positive definite, and invertible. The direction $M_{k}^{-1}\nabla\Phi(\xi_{k})$, satisfies (5.5). Then Algorithm˜1 terminates after finitely many steps. Due to Lemma˜4.1 the returned element $d_{k}$ is a descent direction to the Fréchet differentiable function $\Phi$. Under this condition, the Armijo linesearch terminates after finitely many steps.
The termination criterion is motivated by the following bound, see also [HinzeVierling2012oms, Lemma 4.3 (iii)].
@@
which proves the claim using the Cauchy-Schwarz inequality.
@@
Let $(\xi_{k})$ be a sequence of iterates of Algorithm˜2 together with the search directions $(d_{k})$ satisfying (5.5). Then it holds
(5.7) $\langle d_{k},\nabla\Phi(\xi_{k})\rangle\leq-\frac{1}{1+c}\|\nabla\Phi(\xi_{k})\|_{Y}^{2},$
where $c$ is from ˜1.
Proof 5.10.
Lemma˜4.1 implies $\langle d_{k},-\nabla\Phi(\xi_{k})\rangle\geq\frac{1}{\|M_{k}\|_{\mathcal{L}(Y,Y)}}\|\nabla\Phi(\xi_{k})\|_{Y}^{2}$, and the claim follows with Corollary˜5.4.
@@
Let $(\xi_{k})$ be a sequence of iterates of Algorithm˜2 together with the step sizes $(t_{k})$ and search directions $(d_{k})$. Then
@@
Proof 5.12.
Due to Lemma˜2.7, $G^{*}$, $\Psi$, and $\Phi$ are bounded from below, hence $(\Phi(\xi_{k}))$ is monotonically decreasing and converging. The claim is now a consequence of (5.6).
@@
Using Corollary˜5.1, we will prove a lower bound of $t_{k}$ next. There, we will use the global Lipschitz continuity of $\nabla\Phi$. If we would know that $(\xi_{k})$ has strongly converging subsequences then local Lipschitz continuity of $\nabla\Phi$ would be enough.
@@
Let $(\xi_{k})$ be a sequence of iterates of Algorithm˜2 together with the step sizes $(t_{k})$ and search directions $(d_{k})$. Then there is a constant $\sigma^{\prime}>0$ such that $t_{k}<1$ implies
@@
If $t_{k}<1$ satisfies the Armijo condition (5.6) then $l_{k}\geq 1$, and $\beta^{l_{k}-1}=t_{k}/\beta$ violates the Armijo condition (5.6). Consequently, $t_{k}/\beta$ is larger than the quantity on the right-hand side in the above inequality, which implies the claim.